The probability of raining on Sunday is 0.07. If today is Sunday then find the probability of rain today.
Answer #1

P(Rain and Sunday) = 0.07

P(Sunday) = 1 / 7 (Since 7 days in a week)

We have to calculate P(rain | Sunday) = ?

Using conditional probability formula,

P(rain | Sunday) = P(Rain and Sunday) / P(Sunday)

= 0.07 / (1 / 7)

= 0.07 * 7

= 0.49